# The quark gap equation in light-cone gauge

### v1

@@ IV Gauge dependence of the quark propagator
We compare numerical solutions of the quark’s dressing functions for both, the vertex in Eq. (19) and the bare vertex, that is the solutions of Eqs. (38), (39) and (40) in the so-called rainbow approximation and their counterpart using the WFGTI-vertex. We first discuss the case of the bare vertex for which we choose $D\omega=(2.912\;\mathrm{GeV})^{3}$ and $\omega=0.4$ GeV in Eq. (37). The dressing functions $A(p,n)$ and $B(p,n)$ are presented in Fig. 1 from which we infer that their functional behavior is reminiscent of that found in covariant gauges Lessa:2022wqc . In Fig. 2 one observes the typical rapid increase of the mass function at a hadronic scale of about 1 GeV. Unlike the two other dressing functions $C(p,n)$ is complex valued, as illustrated in Fig. 3. The imaginary part of $C(p,n)$ tends to a finite small value at very low momenta and falls off rapidly after $p^{2}\gtrsim 1\;\mathrm{GeV}^{2}$, whereas the real part remains zero for all $p^{2}$ values. As we shall shortly see, we deal with a special case and the real part of $C(p,n)$ is generally non zero.
@@
Figure 5: The dependence of the quark-mass function $M(p,n)=B(p,n)/A(p,n)$ on the orientation of the light-like vector $n$, see Eq. (47).
Turning to the case of the dressed vertex that satisfies the WFGTI (20), the integral equations for $A(p,n)$, $B(p,n)$ and $C(p,n)$ are more lengthy and the convergence of the numerical iterative procedure is considerably slower. So as to produce a constituent mass of comparable size than the one found with the RL truncation, we readjust the gluon interaction (37): $D\omega=(0.917\;\mathrm{GeV})^{3}$ and $\omega=0.4$ GeV. The wave renormalization and mass functions for this vertex are compared to the ones of the RL vertex in Fig. 1. Both, $A(p,n)$ and $B(p,n)$, are significantly suppressed in this case, yet the mass functions are qualitatively and quantitatively alike. Moreover, Fig. 3 shows that $C(p,n)$ is again a purely imaginary, monotonically decreasing function of $p^{2}$ which tends to a finite negative value in the small-momentum limit. It is worthwhile mentioning that for either vertex our results corroborate the analytic expression found in a one-loop calculation in light-cone gauge Mirja2020 , which shows that $C(p,n)$ is imaginary and decreases rapidly at large momenta.
@@
For convenience, the solutions presented in Figs. 1, 2 and 3 were obtained in the rest frame of the quark, $p=(\mathbf{0},|p|)$. In order to verify the dependence on the light-like vector $n=(0,0,1,i)$, we solve the DSE for an arbitrary momentum $p$, where the scalar product of both vectors is defined as,
@@
with angles $y_{p}=\cos\theta_{p}$ and $z_{p}=\cos\psi_{p}$. This allows us to investigate how the three dressing functions of the quark propagator explicitly depend on the relative orientation of the four-vectors $n$ and $p$. A rigorous calculation requires that $A(k,n)$, $B(k,n)$ and $C(k,n)$ in the kernels of Eqs. (38), (39) and (40) be functions of the three variables $p$, $y_{p}$ and $z_{p}$. Therefore, in solving iteratively this system of coupled integral equations, the values of $A(p,n)$, $B(p,n)$ and $C(p,n)$ must be obtained on a three-dimensional mesh for discrete values $p^{i}$, $y_{p}^{k}$ and $z_{p}^{l}$ in a first iteration and then fed back into the integral equations.
For simplicity’s sake, we here only consider the angular dependences of the explicit $n\cdot p$ terms in Eqs. (38), (39) and (40), and neglect it in the scalar functions which are simplify taken as: $A(k)$, $B(k)$ and $C(k)$. This suffices to reveal their dependence on the orientation of the light-like vector $n$. The behavior of the wave-renormalization and mass functions as functions of $y_{p}$ and $z_{p}$ in the RL-truncation is illustrated in Figs. 4 and 5 for a sample of representative angles. We observe a rapid increase of $A(p,k)$ and conversely an important suppression of $B(p,k)$ when the angles depart from their initial values $y_{p}=z_{p}=1\,\Rightarrow\,n\cdot p=i|p|$. Consequently, $M(p,k)$ is also suppressed. Note that with the present simplification of the integral equations, $A(p,n)$ and $B(p,n)$ are identical for pairs of angles, $y_{p},z_{p}$ and $y_{p}^{\prime}=z_{p},z_{p}^{\prime}=y_{p}$. Similarly, we note that for fixed values of $y_{p}$ these functions do not exhibit any variations but $C(p,n)$ rapidly falls off with increasing steepness as a function of $z_{p}$.
Figure 6: Angular gauge dependence of the real and imaginary parts of the dressing function $C(p,n)$ calculated with a bare vertex.
The $n$-dependent variation of the complex dressing function $C(p,n)$ for the same set of angles is depicted in Fig. 6. Clearly, when $y_{p}\neq 1$ and $z_{p}\neq 1$ the real part of $C(p,n)$ is non-zero for $p^{2}>0$ GeV2 and its functional behavior mirrors the one of the imaginary part, decreasing rapidly with an $n$-dependent steepness for momenta larger than $|p|\gtrsim 1\;\mathrm{GeV}$.
@@ V Concluding remarks
Motivated by phenomenological considerations, namely the intricacy of treating a Wilson line between two quark fields without resorting to expansions or simply neglecting the operator in the definition of distribution functions, we made a first step towards calculating the meson’s wave function in light-cone gauge. To that end, we have solved for the first time the DSE of a quark employing a model dressing function of the gluon propagator in light-cone gauge. We showed that the ML prescription is moot in solving the relevant DSE and that the dual vector $n^{*}$ is not needed.
@@
As a consequence, the solutions of the nonperturbative quark propagator contain three scalar functions, two of them playing the usual role of the mass and wave-renormalization functions, while a third dressing function is complex. All three function depend on the relative orientation of the quark momentum $p$ and the light-like vector $n$. As such, the breaking of covariance in light-cone gauge is expressed by a geometric gauge dependence. The angular dependence of $A(p,n)$, $B(p,n)$ and $C(p,n)$ is found to be significant and in particular the mass function becomes suppressed when moving away from the quark’s rest frame. Naturally, a rigorous discussion of the quark gap equation in light-cone gauge requires the solution the gluon in the same DSE framework with its full dependence on the vector $n$, which may or may not compensate this suppression of the mass function.
@@
Finally, since in light-cone gauge the gluon decouples from the ghost, we derive a dressed quark-gluon vertex that satisfies its WFGTI. With a readjustment of the gluon-dressing function we find that the dressing functions $A(p,n)$, $B(p,n)$ are suppressed in comparison with the same functions in the bare-vertex truncation of the DSE, nonetheless the resulting mass function is qualitatively and quantitatively equivalent. Since the convergence of the iterative procedure using Newton’s method is very slow in case of the WFGTI vertex for arbitrary values of $n\cdot p$, we refrained from studying the corresponding angular dependence of the dressing functions. Future improvements should implement the full angular dependence in the integral equations for $A(p,n)$, $B(p,n)$ and $C(p,n)$ and improve their numerical treatment.
@@
With regard to the gauge independence of physical observables, only a symmetry preserving Bethe-Salpeter kernel for quark-antiquark states in the same light-cone gauge framework can offer a sensible path. This difficult task has been pursued in Landau gauge and should be seriously investigated in light-cone gauge in order to be able to calculate mesonic distribution amplitudes and parton distribution functions.
@@ Acknowledgements.
We would like to express our gratitude to Peter Tandy who suggested the study of the quark in light-cone gauge and who has been of great help to interprete our findings. We also acknowledge stimulating discussions with Gastao Krein. B. E. and R. C. S participate in the Brazilian network project INCT-Física Nuclear e Aplicações, grant no. 464898/2014-5. This work was supported by the São Paulo Research Foundation (FAPESP), grant no. 2023/00195-8 and by the National Council for Scientific and Technological Development (CNPq), grant no. 409032/2023-9.

### v2

@@ IV Gauge dependence of the quark propagator
We compare numerical solutions of the quark’s dressing functions for both, the bare vertex and the vertex in Eq. (19), that is the solutions of Eqs. (38), (39) and (40) in the rainbow approximation and their counterpart using the WFGTI-vertex. In both cases we use the values, $m_{g}=0.3$ GeV and $\omega=0.15$ GeV, in Eq. (36), while $b=2.0\times 10^{-4}$ for the bare vertex and $b=9.0\times 10^{-3}$ for the WFGTI-vertex. The dressing functions $A(p,n)$ and $B(p,n)$ are presented in Fig. 1 in the quark’s rest frame, $p=(\mathbf{0},|p|)$, from which we infer that their functional behavior is reminiscent of that found in covariant gauges Lessa:2022wqc . In Fig. 2 one observes the typical rapid increase of the mass function at a hadronic scale of about 1 GeV and $M(0)\approx 210$ MeV when the bare vertex is employed. Decreasing the value of $b$ increases the mass function, though the solutions are unstable. We choose the parameters, $m_{g}$ and $\omega$, to be the same in both cases, however, slight modifications of them results in a constituent mass above 300 MeV even when solving the DSE with the bare vertex. Unlike the two other dressing functions $C(p,n)$ is complex valued, as illustrated in Fig. 3. Both the real and imaginary parts of $C(p=\mu,n)$ are zero due to our renormalization prescription, and one merely observes insignificant oscillations of these function consistent with $C(p,n)\approx 0$ in the entire range of $p^{2}$ considered.
@@
Figure 5: The dependence of the quark-mass function $M(p,n)=B(p,n)/A(p,n)$ on the quark momentum $p$ in light-cone direction, see Eq. (47).
Turning to the case of the dressed vertex that satisfies the WFGTI (20), the integral equations for $A(p,n)$, $B(p,n)$ and $C(p,n)$ are more lengthy and the convergence of the numerical iterative procedure is considerably slower. So as to produce a constituent mass that is comparable with that of the bare vertex, we readjust the gluon dressing with $b=9.0\times 10^{-3}$, as already mentioned. The $A(p,n)$ and $B(p,n)$ functions are enhanced compared with those of the bare vertex in Fig. 1. While the mass functions are qualitatively alike below the renormalization point, $\mu=2$ GeV, with a steeper rise in case of the WFGTI vertex, the behavior is different at larger momenta: the mass function continuously decreases when a bare vertex is employed, whereas it saturates at about 20 MeV with the WFGTI vertex. Moreover, in Fig. 3 we observe again that $C(p,n)\approx 0$.
@@
For convenience, the solutions presented in Figs. 1, 2 and 3 were obtained in the rest frame of the quark. In order to verify the dependence on the orientation of the quark momentum with respect to the light-like vector $n=(0,0,1,i)$, we solve the DSE for an arbitrary momentum $p$. The scalar product of both vectors is then defined as,
@@
with angles $y_{p}=\cos\theta_{p}$ and $z_{p}=\cos\psi_{p}$. This allows us to investigate how the three dressing functions of the quark propagator explicitly depend on the relative orientation of the four-vectors $n$ and $p$. A rigorous calculation requires that $A(k,n)$, $B(k,n)$ and $C(k,n)$ in the kernels of Eqs. (38), (39) and (40) be functions of the three variables $p$, $y_{p}$ and $z_{p}$. Therefore, in solving iteratively this system of coupled integral equations, the values of $A(p,n)$, $B(p,n)$ and $C(p,n)$ must be obtained on a three-dimensional mesh for discrete values $p^{i}$, $y_{p}^{k}$ and $z_{p}^{l}$ in a first iteration and then fed back into the integral equations. As this is an exploratory study, we limit ourselves to the angular dependences of the explicit $n\cdot p$ terms in Eqs. (38), (39) and (40). For our purposes, this suffices to demonstrate the dependence on the orientation of the quark momentum relative to the light-like vector $n$. We note that the denominator of the third term in Eq. (40) possesses a characteristic $(n\cdot q)^{-1}$ pole when $z_{k}=+1$ and $y_{k}=0$, where the angles are between the vectors $n$ and $k$. In treating the integral numerically, we limit ourselves to $1\leq z_{k}\leq+1-\varepsilon$ with $\varepsilon\approx 10^{-3}$ and use the Cauchy principal value about the value $y_{k}=0$.
The behavior of the wave-renormalization and mass functions as functions of $y_{p}$ and $z_{p}$ in the bare-vertex truncation is illustrated in Figs. 4 and 5 for a sample of representative angles. We observe an increase of $A(p,k)$ in the infrared domain $p^{2}\lesssim 1\;\text{GeV}^{2}$ while the variation in $B(p,k)$ is less marked when the angles depart from their initial values $y_{p}=z_{p}=1\,\Rightarrow\,n\cdot p=i|p|$. Consequently, $M(p,k)$ is also suppressed for decreasing $z_{p}$. Note that $A(p,n)$ and $B(p,n)$ are identical for pairs of angles, $y_{p},z_{p}$ and $y_{p}^{\prime}=z_{p},z_{p}^{\prime}=y_{p}$, considering only the explicit $n\cdot p$ terms.
Figure 6: Light-cone gauge dependence of the real and imaginary parts of the dressing function $C(p,n)$ calculated with a bare vertex.
The $n$-dependent variation of the complex dressing function $C(p,n)$ for the same set of angles is depicted in Fig. 6. It appears that when $y_{p}\neq 1$ and $z_{p}\neq 1$ the real and imaginary parts of $C(p,n)$ slightly oscillate about the renormalization point $\mu=2$ GeV. We verified that these are not simply numerical fluctuations in the integration using both a deterministic cubature rule and a Monte Carlo method, namely the CUHRE and VEGAS algorithms of the CUBA library Hahn:2004fe . At this point, not including the full $n$-dependence in the integral (40), it is difficult to decide whether $C(p,n)=0$ for all momenta and possible angles.
The convergence of the quark DSE with the vertex ansatz in Eq. (19) is considerably slower for the values $z_{p}<1$ and $y_{p}<1$ and solutions become unstable. We therefore abstain from presenting variations of $y_{p}$ and $z_{p}$ for this case.
@@ V Concluding remarks
Motivated by phenomenological considerations, namely the intricacy of treating a Wilson line between two quark fields in nonperturbative functional approaches without resorting to expansions or simply neglecting the operator in the definition of distribution functions, we made a first step towards a meson’s wave function in light-cone gauge. To that end, we solved for the first time the DSE of a quark in light-cone gauge employing Cornwall’s gluon propagator Cornwall:1981zr . We showed that the ML prescription is moot in solving the relevant DSE and that the dual vector $n^{*}$ is not needed.
@@
As a consequence, the numerical solutions of the nonperturbative quark propagator contain three scalar functions, two of them playing the usual role of the mass and wave-renormalization functions, while a third dressing function is complex. The dependence of $A(p,n)$ and $B(p,n)$ on the relative orientation of the quark momentum $p$ in light-cone direction is found to be only significant below 1 GeV, and the mass function is suppressed when moving away from the quark’s rest frame. On the other hand, the real and imaginary parts of $C(p,n)$ are consistent with zero in the quark’s rest frame, whereas in a moving frame $C(p,n)$ oscillates about the renormalization point. In other words, the breaking of covariance in light-cone gauge is manifest in the scalar functions.
Naturally, a rigorous discussion of the quark gap equation in light-cone gauge requires the self-consistent solutions of the quark and the gluon in the same DSE framework and with their full dependence on the vector $n$. This may or may not compensate the suppression of the mass function. According to Cornwall, the dressing function of the gluon propagator is a covariant scalar function that only depends on its momentum Cornwall:1981zr . One may extend this reasoning to the quark propagator and insist that fully and correctly renormalized scalar Green functions of one momentum should be covariant. Leibbrandt’s regularization prescription to handle $1/n\cdot q$ singularities in light-cone gauge, on the other hand, leads to expressions for $A(p,n)$, $B(p,n)$ and $C(p,n)$ that do depend on the light-like vector, $n$, and in Ref. Mirja2020 his calculation was extended to include divergent terms renormalized in the $\overline{\text{MS}}$ scheme. Nevertheless, at the level of individual Feynman diagrams the singularities can lead to gauge-dependent terms proportional to the four-vector $n$ in self-energy corrections. Ultimately, these dependences might cancel. Whichever it is, the quark propagator is not an observable and gauge dependent. Therefore, even if $C(p,n)\equiv 0$ was a true statement, the gauge dependence should manifest itself in the wave-function renormalization and mass function. At this point, we cannot answer the question of how the quark propagator depends on the light-like vector.
@@
Since in light-cone gauge the gluon decouples from the ghost, we also derived a dressed quark-gluon vertex that satisfies its WFGTI. With a readjustment of the gluon-dressing function we find that the dressing functions $A(p,n)$, $B(p,n)$ are suppressed in comparison with the same functions in the bare-vertex truncation of the DSE, nonetheless the resulting mass function is qualitatively and quantitatively equivalent. Since the convergence of the iterative procedure using Newton’s method is very slow in case of the WFGTI vertex for arbitrary values of $n\cdot p$, we refrained from studying the mass function’s dependence on it in this case. We remind that future improvements should implement the full angular dependence in the integral equations for $A(p,n)$, $B(p,n)$ and $C(p,n)$.
@@
With regard to the gauge independence of physical observables, only a symmetry preserving Bethe-Salpeter kernel for quark-antiquark states in the same light-cone gauge framework can offer a sensible path. This difficult task has been pursued in Landau gauge and ought to be investigated in light-cone gauge in order to be able to calculate mesonic distribution amplitudes and parton distribution functions.
@@ Acknowledgements.
We would like to express our gratitude to Peter Tandy who suggested the study of the nonperturbative quark propagator in light-cone gauge and who has been of great help to interprete our findings. We also appreciated stimulating discussions with Arlene Cristina Aguilar and Gastão Krein. B. E. and R. C. S participate in the Brazilian network project INCT-Física Nuclear e Aplicações. This work was supported by the São Paulo Research Foundation (FAPESP), grant no. 2023/00195-8 and by the National Council for Scientific and Technological Development (CNPq), grant no. 409032/2023-9. R. C. S is supported by CAPES fellowship, grant no. 88887.086709/2024-00.